# Relational Conformal Prediction for Correlated Time Series

### v1

@@ 5.1 Time series forecasting benchmarks
We consider the following datasets, each coming from a different application domain: METR-LA from the traffic forecasting literature [65]; a collection of air quality measurements from different Chinese cities (AQI[66]; a collection of energy consumption profiles acquired from smart meters within the CER smart metering project (CER-E[67, 68]. We follow the preprocessing steps of previous works [65, 69, 37] and adopt $40\%/40\%/20\%$ splits for training, calibration, and testing, respectively. For each dataset, we first train $3$ different baseline models: a simple RNN with gated recurrent unit (GRU) cells [70], a decoder-only Transformer [71], a simple TTS STGNN by following the template in Sec. 2.4. The latter uses a pre-defined graph that models the dependencies across the time series. After training, we evaluate each baseline on the calibration set and save the associated residuals, which are then used as input to the different CP methods. More details on the datasets and base models are provided in App. C.
@@
We compared CoRel against the following methods: 1) SCP, the standard split CP; 2) SeqCP, where, analogously to Xu and Xie [18], we compute empirical quantiles using only the most recent $K$ residuals at each time step; 3) NexCP [16], which computes empirical quantiles by assigning exponentially decaying weights to past residuals; 4) SCPI [19], which estimates the residuals’ quantile function from the last few steps of each time series with a random forest; 5) HopCPT which reweights past residuals by learning attention scores with a Modern Hopfield Network [72]. We include in our comparison a model called CoRNN, where we use the same architecture as CoRel but remove the message-passing layers. Except for HopCPT which uses a custom procedure [21], model selection is performed on a validation set by optimizing the Winkler score.
@@
W/o adapt. -3.46$\pm$ 0.16 1.91$\pm$ 0.01 3.83$\pm$ 0.05 -3.07$\pm$ 0.31 1.95$\pm$ 0.03 3.82$\pm$ 0.02 -5.15$\pm$ 0.44 1.79$\pm$ 0.03 3.79$\pm$ 0.04
W/ adapt. -2.65$\pm$ 0.12 1.93$\pm$ 0.01 3.61$\pm$ 0.05 -2.39$\pm$ 0.37 1.96$\pm$ 0.01 3.64$\pm$ 0.02 -3.13$\pm$ 0.28 1.85$\pm$ 0.01 3.59$\pm$ 0.03
Results
Tab. 1 reports the results across the datasets and the base prediction models. The first observation is that CoRel outperforms the competitors in terms of Winkler score in almost all cases. We observed a few exceptions only when the baseline is itself an STGNN, as it is already expected to take care of modeling spatiotemporal dependencies. However, note that the STGNN base model has access to a pre-defined graph, which is not always available in practical applications. In terms of coverage, CoRel achieves good results, with the exception of some cases in the CER-E dataset. However, we note that our model selection prioritized the Winkler score which emphasizes the width of the prediction bands besides the coverage. CoRNN, the simplified version of our approach, obtains good overall performance and is competitive against the state-of-the-art but, despite providing good coverage, it is outperformed by CoRel in most scenarios in terms of Winkler score. Among the competitors, HopCPT provides competitive results in METR-LA and CER-E but with a larger coverage gap. Except for SeqCP, the other baselines obtain good coverage in most settings at the expense of drastically wider PIs.
Table 3: Performance on GPVAR.
GPVAR-RNN GPVAR-STGNN
Models $\Delta$Cov Winkler $\Delta$Cov Winkler
SCP -0.023 2.14 -0.019 1.66
CoRNN -0.1$\pm$0.2 2.04$\pm$0.00 0.0$\pm$0.0 1.66$\pm$0.00
CoRel 0.0$\pm$0.1 1.67$\pm$0.00 0.1$\pm$0.1 1.66$\pm$0.00
w/ true ${\bm{A}}$0.1$\pm$0.1 1.66$\pm$0.00 0.0$\pm$0.2 1.66$\pm$0.00
5.2 Controlled environment
We evaluated the behavior of CoRel in a controlled environment by simulating a diffusion process on a graph. In particular, the experiment relied on the GPVAR benchmark introduced by Zambon and Alippi [12], with a setup analogous to [37]. Data were generated recursively from the auto-regressive polynomial graph filter:
${\bm{H}}_{t}$ $=\sum_{l=1}^{L}\sum_{q=1}^{Q}\Theta_{q,l}{\bm{A}}^{l-1}{\bm{X}}_{t-q},$
${\bm{X}}_{t+1}$ $={\bm{a}}\odot\text{tanh}\left({\bm{H}}_{t}\right)+{\bm{b}}\odot\text{tanh}\left({\bm{X}}_{t-1}\right)+\eta_{t},$ (19)
where parameters $\bm{\Theta}\in{\mathbb{R}}^{Q\times L}$, ${\bm{a}}\in{\mathbb{R}}^{N}$, ${\bm{b}}\in{\mathbb{R}}^{N}$ are kept fixed across nodes and $\eta_{t}\sim\mathcal{N}(\bm{0},\sigma^{2}{\mathbb{I}})$. We ran the simulation on a graph with $60$ nodes and with a topology analogous to previous works [37]; more details are provided in App. C. We use both a RNN and STGNN as base point predictors. As shown by Zambon and Alippi [12], STGNNs can obtain a forecasting accuracy near the theoretical optimum in this dataset, which results in i.i.d. residuals. As such, we would expect standard SCP to be sufficient in this scenario. The objective of this experiment is to show that CoRel is effectively able to capture and leverage existing spatiotemporal dependencies.
Results
We compare CoRel against standard SCP and the CoRNN variant. Moreover, we also compare performance against CoRel with access to the true graph used to generate the data. Results are shown in Tab. 3. When using an RNN as a point predictor (GPVAR-RNN) CoRel significantly outperforms both standard SCP and CoRNN. Furthermore, CoRel achieves results that closely match those obtained with direct access to the ground truth graph, showing the effectiveness of the proposed architecture in capturing latent relational dependencies. Finally, results obtained by using an STGNN as baseline (GPVAR-STGNN) show (as expected) that the standard SCP is sufficient when the point-predictor captures all the relevant dependencies.
5.3 Adaptability
In this experiment, we evaluate how effectively the adaptation technique proposed in Sec. 3.4 provides accurate PIs over time. We focused on the CER-E dataset where the calibration set does not cover a full year, which likely introduces a shift at test time given the seasonality of energy consumption. This behavior is also confirmed by the results in Sec. 5.1.
We used CoRel with a fixed hyperparameter configuration across scenarios and trained on the calibration set. We then split the test set in $K=6$ folds and then iteratively evaluated and fine-tuned the model on each fold. Results, reported in Tab. 2, show that the proposed adaptation schemes improve performance by reducing the coverage gap and improving the accuracy of the PI.
@@ 6 Conclusion
Future works
We believe that CoRel constitutes an important step toward effective spatiotemporal CP methods. There are several directions for future works to explore, such as applying the framework to heterogenous time series. Future work should also further focus on mitigating issues coming from learning in non-stationary environments.
@@
For $\hat{C}^{\alpha}_{t}$ in both Eq. 14 and Eq. 16 we have by construction that
$G^{c}_{\psi}({\bm{X}}_{t+H}\in\hat{C}^{\alpha}_{t})=1-\alpha.$ (22)
@@
$P^{c}_{t+H}\left({\bm{X}}_{t+H}\in\hat{C}^{\alpha}_{t}(\widehat{\bm{X}}_{t+h})\right)\geq 1-\alpha-{TV}\left(P^{c}_{\psi},P^{c}_{t+H}\right).$
@@ Appendix B Hardware and software platforms
Benchmarks have been developed with Python [73] and the following open-source libraries:
  • Numpy [74];
  • PyTorch [75];
  • PyTorch Lightning [76];
  • PyTorch Geometric [77];
  • Torch Spatiotemporal [78].
Experiments were conducted on a server equipped with AMD EPYC 7513 CPUs and NVIDIA RTX A5000 GPUs. The code for reproducing the computational experiments will be open-sourced upon publication. For the HopCPT, SCPI, NexCP, and SeqCP baselines we use the open-source implementation made available by Auer et al. [21] [^1] https://github.com/ml-jku/HopCPT.
@@ Appendix C Datasets
We considered several datasets from different application domains, real-world scenarios and a simulated controlled environment. For the real world datasets we followed the pre-processing steps adopted by Cini et al. [37].
@@
The AQI [66] dataset consists of 8,760 hourly measurements of pollutant PM2.5 from 437 monitoring stations in China. We use window of $W=24$ time steps and predict the $3$-time-step-ahead observations. For the STGNN base model, we build an adjacency matrix using a thresholded Gaussian kernel computed from pairwise geographic distances [79].
@@
We considered the METR-LA [65] traffic forecasting dataset, consisting of 34,272 timesteps of measurements from 207 loop detectors sampled at 5-minute intervals in the Los Angeles County Highways. We use window of $12$ time steps and predict the $12$-time-step-ahead observations. For the STGNN base model, we followed previous works [69] and build an adjacency matrix using a thresholded Gaussian kernel applied to geographic distances.
@@
We selected the CER-E dataset [67, 68], comprising 25,728 timesteps of energy consumption readings aggregated at 30-minute intervals from 485 smart meters monitoring small and medium-sized enterprises. We use window of $48$ time steps and predict the $5$-time-step-ahead observations. For the STGNN base model, we built the adjacency matrix by extracting a 10-nearest neighbor graph from week-wise correntropy similarities between time series, following previous work [68].
@@
For the GPVAR dataset, we generate synthetic data with 40,000 timesteps over an undirected network of 60 nodes connected in a community graph structure by following the system model in Sec. 5.2 [12]. The parameters of the spatiotemporal process are set as
$\Theta=\begin{bmatrix}2.5&-2.0&-0.5\\ 1.0&3.0&0.0\end{bmatrix},\quad a=b=0.5.$ (23)
@@
We trained three base models (point predictors) for each dataset: a RNN with GRU cells (1 layer with hidden size 32), a decoder-only Transformer (hidden size 32, feed-forward size 64, 2 attention heads, 3 layers, dropout 0.1), and a STGNN following the template in Sec. 2.4 (hidden size 32, node embedding size 16, 1 layer GRU, 2 message-passing layers). All models were trained by minimizing the MAE loss using the Adam optimizer for 200 epochs with batch size 32, using 40% of the data for training, 40% for calibration, and 20% for testing. Input features were standardized across the spatial and temporal dimensions. For the STGNN, we used the pre-defined graph structure provided with each dataset.
@@ Appendix D Additional details on CoRel implementation
As discussed in Sec. 3 and Sec. 5, we implemented CoRel as TTS model with a single-layer GRU followed by $2$ message passing layers. As a readout, we used an MLP mapping the learned representations to prediction of $39$ equally spaced quantiles.
The graph learning module was implemented as described in Sec. 3.2 by parametrizing ${\bm{\Phi}}$ with a matrix of $N\times N$ learnable parameters. To allow for sampling less neighbors than the specified neighborhood size $K$, we followed previous works [40] and modified the sampling procedure by introducing a set of dummy nodes then discarded from the sampled graph before message passing. We use dummy nodes in the GPVAR experiment only. To allow for sparse message-passing operations we use a straight-through gradient estimator [45] and backpropagate gradient only through the sampled edges for each node plus $10\%$ of the remaining ones. In the GPVAR experiment we simply propagate gradients w.r.t. the entire adjacency matrix.
Besides residuals, we use as additional covariates datetime encodings whenever available plus the value of the target time series w.r.t. the time steps in the input window.

### v2

@@ 5.1 Time Series Forecasting Benchmarks
We consider the following datasets, each coming from a different application domain: METR-LA from the traffic forecasting literature (Li et al., 2018); a collection of air quality measurements from different Chinese cities (AQI(Zheng et al., 2015); a collection of energy consumption profiles acquired from smart meters within the CER smart metering project (CER-E(Commission for Energy Regulation, 2016; Cini et al., 2022). We follow the preprocessing steps of previous works (Li et al., 2018; Wu et al., 2019; Cini et al., 2023b) and adopt $40\%/40\%/20\%$ splits for training, calibration, and testing, respectively. For each dataset, we first train $3$ different baseline models: a simple RNN with gated recurrent unit (GRU) cells (Cho et al., 2014), a decoder-only Transformer (Vaswani et al., 2017), and a simple TTS STGNN obtained by following the template in Sec. 2.4. The latter uses a pre-defined graph that models the dependencies across the time series. After training, we evaluate each baseline on the calibration set and save the associated residuals, which are then used as input to the different CP methods. More details on the datasets and base models are provided in App. C.
@@
We compared CoRel against the following baselines: 1) SCP, the standard split CP; 2) SeqCP, where, analogously to Xu and Xie (2023a), we compute empirical quantiles using only the most recent $K$ residuals at each time step; 3) NexCP (Barber et al., 2023), which computes empirical quantiles by assigning exponentially decaying weights to past residuals; 4) SPCI (Xu and Xie, 2023b), which estimates the residuals’ quantile function from the last few steps of each time series with a random forest; 5) HopCPT which reweights past residuals by learning attention scores with a Modern Hopfield Network (Ramsauer et al., 2021). These baselines are representative of the current state-of-the-art in CP for time series forecasting; in particular, SPCI (based on quantile regression) and HopCPT (based on reweighting) are representative of the main recent paradigms. Note that a comparison with non-post-hoc methods would be problematic, as results heavily depend on the base predictors being used. We also include in our comparison a model called CoRNN, where we use the same architecture as CoRel but remove message-passing layers and node embeddings. CoRNN is an ablation of the introduced designs. Except for HopCPT, which uses a custom procedure (Auer et al., 2023), model selection is performed on a validation set by optimizing the Winkler score.
Results
Tab. 1 reports the results across the datasets and the base prediction models. The first observation is that CoRel outperforms the competitors in terms of Winkler score in almost all cases. We observed a few exceptions only when the baseline is itself an STGNN, as it is already expected to take care of modeling spatiotemporal dependencies. However, note that the STGNN base model has access to a pre-defined graph, which is not always available in practical applications. In terms of coverage, CoRel achieves good results, with the exception of some cases in the CER-E dataset. However, we note that our model selection prioritized the Winkler score which emphasizes the width of the prediction bands besides the coverage. CoRNN, the simplified version of our approach, obtains good overall performance and is competitive against the state-of-the-art but, despite providing good coverage, it is outperformed by CoRel in most scenarios in terms of Winkler score. Among the competitors, HopCPT provides competitive results in METR-LA and CER-E but with a larger coverage gap. Except for SeqCP, the other baselines obtain good coverage in most settings at the expense of drastically wider PIs. Regarding computational scalability, note that CoRel shares most of the learnable parameters among time series and that its training can be efficiently parallelized on a GPU. Furthermore, besides learnable node embeddings ${\bm{V}}$, CoRel relies only on a short window of the most recent observations as an input at each time step. Conversely, SPCI requires training a different model for each time series, and HopCPT requires computing attention scores w.r.t. the entire calibration set at each time step.
5.2 Controlled Environment
Table 2: Performance on GPVAR.
GPVAR-RNN GPVAR-STGNN
Models $\Delta$Cov PI-Width Winkler $\Delta$Cov PI-Width Winkler
SCP -0.02 1.67 2.14 -0.02 1.32 1.66
CoRNN -0.1$\pm$0.2 1.63$\pm$0.01 2.04$\pm$0.00 0.0$\pm$0.0 1.32$\pm$0.00 1.66$\pm$0.00
CoRel 0.0$\pm$0.1 1.33$\pm$0.00 1.67$\pm$0.00 0.1$\pm$0.1 1.32$\pm$0.01 1.66$\pm$0.00
w/ true ${\bm{A}}$0.1$\pm$0.1 1.32$\pm$0.00 1.66$\pm$0.00 0.0$\pm$0.2 1.32$\pm$0.01 1.66$\pm$0.00
We evaluated the behavior of CoRel in a controlled environment by simulating a diffusion process on a graph. In particular, the experiment relied on the GPVAR benchmark introduced by Zambon and Alippi (2022), with a setup analogous to (Cini et al., 2023b). Data were generated recursively from the auto-regressive polynomial graph filter:
${\bm{H}}_{t}$ $=\sum_{l=1}^{L}\sum_{q=1}^{Q}\Theta_{q,l}{\bm{A}}^{l-1}{\bm{X}}_{t-q},$
${\bm{X}}_{t+1}$ $=a\odot\text{tanh}\left({\bm{H}}_{t}\right)+b\odot\text{tanh}\left({\bm{X}}_{t-1}\right)+\eta_{t},$ (19)
where parameters $\bm{\Theta}\in{\mathbb{R}}^{Q\times L}$, $a\in{\mathbb{R}}$, $b\in{\mathbb{R}}$ are kept fixed across nodes and $\eta_{t}\sim\mathcal{N}(\bm{0},\sigma^{2}{\mathbb{I}})$ with $\sigma=0.4$. We ran the simulation on a graph with $60$ nodes and with a topology analogous to previous works (Cini et al., 2023b); more details are provided in App. C. We use both a RNN and STGNN as base point predictors. As shown by Zambon and Alippi (2022), STGNNs can obtain a forecasting accuracy near the theoretical optimum in this dataset, which results in uncorrelated residuals. As such, we would expect standard SCP to be sufficient when using an STGNN as base model. The objective of this experiment is to show that CoRel is effectively able to capture and leverage existing spatiotemporal dependencies.
Results
We compare CoRel against standard SCP and the CoRNN variant. Moreover, we also compare performance against CoRel with access to the true graph used to generate the data. Results are shown in Tab. 2. When using an RNN as a point predictor (GPVAR-RNN) CoRel significantly outperforms both standard SCP and CoRNN. Furthermore, CoRel achieves results that closely match those obtained with direct access to the ground truth graph, which shows the effectiveness of the proposed architecture in capturing latent relational dependencies. Note that, given the injected Gaussian noise with $\sigma=0.4$, the theoretical optimum PI width to obtain (asymptotically) $90\%$ marginal coverage is $1.315$. CoRel achieves essentially perfect coverage with PI width close to the theoretical optimum, while CoRNN requires a substantially higher PI width to obtain similar coverage. Finally, results obtained by using an STGNN as baseline (GPVAR-STGNN) show (as expected) that the standard SCP is sufficient when the point-predictor captures all the relevant dependencies. We provide a visualization of the learned graph in Sec. G.1.
5.3 Adaptation
@@
W/o adapt. -3.44 $\pm$ 0.40 1.84 $\pm$ 0.02 3.71 $\pm$ 0.03 -3.08 $\pm$ 0.15 1.84 $\pm$ 0.02 3.69 $\pm$ 0.04 -4.07 $\pm$ 0.47 1.80 $\pm$ 0.03 3.78 $\pm$ 0.03
W/ adapt. -2.70 $\pm$ 0.27 1.85 $\pm$ 0.01 3.49 $\pm$ 0.02 -2.42 $\pm$ 0.20 1.85 $\pm$ 0.01 3.49 $\pm$ 0.04 -3.06 $\pm$ 0.25 1.81 $\pm$ 0.02 3.58 $\pm$ 0.02
In this experiment, we evaluate how effectively the adaptation technique proposed in Sec. 3.4 provides accurate PIs over time. We focused on the CER-E dataset where the calibration set does not cover a full year, which likely introduces a shift at test time given the seasonality of energy consumption. We used CoRel with a fixed hyperparameter configuration across scenarios and trained on the calibration set. Embeddings are updated every $M$ time steps by running the training procedure on the latest observations and keeping all the parameters frozen except for the embeddings. More details on the hyperparameters are provided in App. B. In practice, this is done by splitting the test set into $K=6$ folds and then iteratively fine-tuning the model on each fold. This procedure simulates a real-world scenario where new data become available over time and are used for fine-tuning. Results, reported in Tab. 3, show that this adaptation scheme improves performance by reducing the coverage gap and providing more accurate PIs.
@@ 6 Conclusion
Limitations and future work
We believe that CoRel constitutes an important step toward effective spatiotemporal CP methods. There are several directions for future work to explore, such as applying the framework to heterogeneous time series. Future work should focus on mitigating challenges arising from learning in non-stationary environments, e.g., by developing methods and coverage guarantees under specific assumptions about the non-stationarity of the data-generating process. Applications of CoRel to very large time series collections and combinations thereof with scalable graph-based architectures (Chiang et al., 2019; Cini et al., 2023c) are an additional direction. Finally, while CoRel accounts for dependencies w.r.t. past observations at correlated time series, it provides separate PIs for each time series. It would then be interesting to extend the framework with components that model the joint probability distribution of the time series.
Impact Statement
This paper presents work whose goal is to advance the field of machine learning and time series forecasting. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Acknowledgments
This work was supported by the Swiss National Science Foundation grant no. 225351 (Relational Deep Learning for Reliable Time Series Forecasting at Scale) and no. 204061 (HORD GNN: Higher-Order Relations and Dynamics in Graph Neural Networks), the Hasler Stiftung project Calibrated Uncertainty Estimation for Spatio-Temporal Data, the UKRI CDT in AI for Healthcare (grant no. P/S023283/1), and the Norwegian Research Council grant no. 345017 (RELAY: Relational Deep Learning for Energy Analytics). FMB wishes to thank Nvidia Corporation for donating some of the GPUs used in this project. AC conducted part of this work while at the University of Oxford.
@@ Acknowledgments
For $\hat{C}^{\alpha}_{i,t}$ in both Eq. 14 and Eq. 16 we have by construction that
$P^{c}_{\psi}\left({\bm{x}}^{i}_{t+H}\in\widehat{C}^{\alpha}_{i,t}\right)=1-\alpha.$ (22)
@@
$P^{c}_{t+H}\left({\bm{x}}^{i}_{t+H}\in\widehat{C}^{\alpha}_{i,t}(\widehat{\bm{X}}_{t+h})\right)\geq 1-\alpha-{TV}\left(P^{c}_{\psi},P^{c}_{t+H}\right).$
@@ Appendix B Hardware and software platforms
Benchmarks have been developed with Python [Van Rossum and Drake, 2009] and the following open-source libraries:
  • Numpy [Harris et al., 2020];
  • PyTorch [Paszke et al., 2019];
  • PyTorch Lightning [Falcon and The PyTorch Lightning team, 2019];
  • PyTorch Geometric [Fey and Lenssen, 2019];
  • Torch Spatiotemporal [Cini and Marisca, 2022].
Experiments were conducted on a server equipped with AMD EPYC 7513 CPUs and NVIDIA RTX A5000 GPUs. The code for reproducing the computational experiments is available at https://github.com/andreacini/corel. For the HopCPT, SPCI, NexCP, and SeqCP baselines we use the open-source implementation made available by Auer et al. [2023] [^1] https://github.com/ml-jku/HopCPT. In our setup for the experiment in Sec. 5.1, considering the METR-LA dataset and the baselines that require fitting a model, training and testing require $\approx 3$ days for SPCI, $\approx 11$ hours for HopCPT, and $\approx 5$ minutes for CoRel.
@@ Appendix C Datasets
We considered several datasets from different application domains, real-world scenarios and a simulated controlled environment. For the real world datasets we followed the pre-processing steps adopted by Cini et al. [2023b].
@@
The AQI [Zheng et al., 2015] dataset consists of 8,760 hourly measurements of pollutant PM2.5 from 437 monitoring stations in China. We use a window of $W=24$ time steps and predict the $3$-time-step-ahead observations. For the STGNN base model, we build an adjacency matrix using a thresholded Gaussian kernel computed from pairwise geographic distances [Shuman et al., 2013].
@@
We considered the METR-LA [Li et al., 2018] traffic forecasting dataset, consisting of 34,272 timesteps of measurements from 207 loop detectors sampled at 5-minute intervals in the Los Angeles County Highways. We use a window of $12$ time steps and predict the $12$-time-step-ahead observations. For the STGNN base model, we followed previous works [Wu et al., 2019] and built an adjacency matrix using a thresholded Gaussian kernel applied to geographic distances.
@@
We selected the CER-E dataset [Commission for Energy Regulation, 2016, Cini et al., 2022], comprising 25,728 timesteps of energy consumption readings aggregated at 30-minute intervals from 485 smart meters monitoring small and medium-sized enterprises. We use a window of $48$ time steps and predict the $5$-time-step-ahead observations. For the STGNN base model, we built the adjacency matrix by extracting a $10$-nearest neighbor graph from week-wise correntropy similarities between time series, following previous work [Cini et al., 2022].
@@
For the GPVAR dataset, we generate synthetic data with 40,000 timesteps over an undirected network of 60 nodes connected in a community graph structure by following the system model in Sec. 5.2 [Zambon and Alippi, 2022]. The parameters of the spatiotemporal process are set as
$\Theta=\begin{bmatrix}2.5&-2.0&-0.5\\ 1.0&3.0&0.0\end{bmatrix},\quad a=b=0.5,\quad\sigma=0.4.$ (23)
@@
We trained three base models (point predictors) for each dataset: a RNN with GRU cells (1 layer with hidden size 32), a decoder-only Transformer (hidden size 32, feed-forward size 64, 2 attention heads, 3 layers, dropout 0.1), and a STGNN following the template in Sec. 2.4 (hidden size 32, node embedding size 16, 1 layer GRU, 2 message-passing layers). All models were trained by minimizing the MAE loss using the Adam optimizer for 200 epochs with batch size 32, using 40% of the data for training, 40% for calibration, and 20% for testing. We also use the first $25\%$ of the calibration data as a validation set for early stopping. Input features were scaled using standard scaling across time series.
@@ Appendix D Additional details on CoRel implementation
Architecture
As discussed in Sec. 3 and Sec. 5, we implemented CoRel as TTS model with a single-layer GRU followed by $2$ message passing layers analogous to those in [Satorras et al., 2022]. As a readout, we used an MLP mapping the learned representations to predictions of $39$ equally spaced quantiles. For the CoRNN baseline, we used the same architecture but removed the message passing layers and node embeddings ${\bm{V}}$ and allowed the GRU to have more than one layer.
Latent graph learning module
The graph learning module was implemented as described in Sec. 3.2 by parametrizing ${\bm{\Phi}}$ with a matrix of $N\times N$ learnable parameters. To allow for sampling less neighbors than the specified neighborhood size $K$, we followed previous works [Cini et al., 2023c] and modified the sampling procedure by introducing a set of dummy nodes then discarded from the sampled graph before message passing. We use dummy nodes in the GPVAR experiment only. To allow for sparse message-passing operations we use a straight-through gradient estimator [Bengio et al., 2013] and backpropagate gradients only through the sampled edges for each node, plus $10\%$ of the remaining ones. In the GPVAR experiment, we simply propagate gradients w.r.t. the entire adjacency matrix.
Covariates
Besides residuals, we use as additional covariates datetime encodings whenever available, plus the value of the target time series w.r.t. the time steps in the input window.